OpenStudy (anonymous):

The sum of the measures of two complementary angles exceeds the difference of their measures by 72°. Find the measure of the smaller angle. A. 18° B. 36° C. 54° D. 58°

OpenStudy (ajprincess):

Let the angles be a and b. a+b=90( sum of complementary angles is 90) a+b-(a-b)=72 90-a+b=72 b-a=72-90 b-a=-18 a-b=18 a+b=90 Does that help?

OpenStudy (anonymous):

This helps me a lot by the way.. :)

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

@waterineyes reported for medal hounding! lol

OpenStudy (anonymous):

he he :p

OpenStudy (anonymous):

**whichever post of aj you see

OpenStudy (anonymous):

now don't tell me you are not typing :P

OpenStudy (anonymous):

good idea! ;)

OpenStudy (anonymous):

well, no mods online, we can spam this question as much as we want :P

OpenStudy (anonymous):

mods can see deleted comments.. -_-

OpenStudy (anonymous):

i would look like a fool talking to myself xD

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

damn!

OpenStudy (anonymous):

Let me tell you a fact., that 4th 'offline' person viewing this post is actually a moderator!

OpenStudy (anonymous):

read my previous comment ;)

OpenStudy (anonymous):

its evening here.

OpenStudy (anonymous):

i know that its evening for all 4 of us

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

now, this is getting lamer :P

OpenStudy (anonymous):

damn i missed your last comment :P

OpenStudy (anonymous):

you cracking this joke since your OS birth.

OpenStudy (anonymous):

that w is of lower caps , and name is in bold

OpenStudy (anonymous):

i never said you are bad, at all

OpenStudy (anonymous):

lets get back to helping.